Select the correct answer: The big O complexity of binary search on an array is: O(1) O(log n)  O(n) O(n^2) other.
If you selected other, write your answer below.
Answer #1
Recursive formula for binary search is 
T(n) = T(n/2) + 1
     = T(n/4) + 1 + 1
     = T(n/8) + 1 + 1 + 1
     ......
     ......
     ......
     = T(n/n) + 1 + .... + 1 + 1 + 1 [log(n) +1 terms]
     = T(1) + 1 + .... + 1 + 1 + 1 [log(n) +1 terms]
     = 1 + 1 + .... + 1 + 1 + 1 [log(n) +1 terms]
     = log(n)
     = O(logn)

O(log n)